brass boiler has a base area of 0.15 m square and thickness 1 cm it  boiled water at a rate of 6 kg per minute when placed on a gas stove estimate the temperature of the part of the flame in contact with the boiler thermal conductivity of Glass is 109 per second per metre per Kelvin heat of vaporization of water is 2256*10^3

Grade:11

1 Answers

Arun
5 years ago
Dear student
 

Base area of the boiler, = 0.15 m2
Thickness of the boiler, l = 1.0 cm = 0.01 m
Boiling rate of water, R = 6.0 kg/min
Mass, m = 6 kg
Time, t = 1 min = 60 s
Thermal conductivity of brass, K = 109 J s –1 m–1 K–1
Heat of vaporisation, L = 2256 × 103 J kg–1
The amount of heat flowing into water through the brass base of the boiler is given by:
θ = KA(T1 – T2t / l    ….(i)

 

where,
T1 = Temperature of the flame in contact with the boiler
T2 = Boiling point of water = 100°C
Heat required for boiling the water:
θ = mL … (ii)
Equating equations (i) and (ii), we get:
∴ mL = KA(T1 – T2t / 
T1 – T2 = mLl / KAt
= 6 × 2256 × 103 × 0.01 / (109 × 0.15 × 60)
= 137.98 o C
Therefore, the temperature of the part of the flame in contact with the boiler is 237.98°C.